GlueX TOF Meeting, October 27, 2011

From GlueXWiki
Jump to: navigation, search

Thursday, October 27, 2011
10:30 am EDT
JLab: CEBAF Center, Room F326


  1. Announcements
  2. Minutes from the last meeting
  3. Mechanical design: Tim, Chuck
  4. PMT base design, rates, currents, amplification Beni
  5. Prototype status: Paul, Sasha
  6. Contract Status: Mark



  1. ESNet: 8542553
  2. EVO: EVO site


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2011-4Q on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Eugene Chudakov, Chuck Hutton, Mark Ito (chair), Simon Taylor, Tim Whitlatch, Beni Zihlmann

Review of Minutes from the Last Meeting

We went over the minutes of the October 13th meeting.

Efficiency with Reduced Gain

Beni summarized his estimates of how low we could go with PMT gain and still preserve good efficiency for minimum ionizing particles. He assumes we want to operate at 20 μA current per tube and are running at 107 photons per second. The would give a tube life of roughly two years. Then on all but the innermost split counters, and a 20 mV discriminator threshold we can maintain good efficiency all along the counter without amplifiers. See his wiki page for details.

Other Items

  • Chuck has been in contact with Ian Winger. Ian will send a new model of the half-width light guide.
  • Mark sent a response about our rates and current to the Hamamatsu engineer who requested information.
  • We still need new milestones for the SOW from Paul.
  • Mark circulated Fernando Barbosa's schematics for the splitters. Sasha checked his splitters against the values listed and found agreement. The origin of the reflections reported last time remains a mystery.
  • Sasha checked least count of his CAEN ADC and got a value of 0.25 pC, as advertised, within about 5%.
  • FSU has not had a look at the 8-stage tubes yet.

Mechanical Design

Chuck reported progress on the support scheme for the modules. See his drawings (added in press) for the latest scheme.

Tim stated that they are pushing forward on the design for holding the steel next.

We discussed how to deal with the possibility of having to replace tubes.

  • Beni will ask Dave Mack about what is done in Hall C and will ask Bogdan W??? about Hall A.
  • We discussed the possibility of removing the entire TOF array for a round of tube replacement.
    • In this case Eugene wondered if the detector can be supported to give horizonal surfaces for the glue joints that are being redone.
  • Choice of glue must be made with replacement in mind.
  • Beni proposed that we use grease to facilitate replacement.
  • We will clearly have to have a mechanical design that allows replacement of tubes.

Prototype Progress

Sasha is now doing measurements with the same splitters on all six signals. It takes about two weeks for one configuration. He is working with the 50/50 split now. He will change to 80/20 next.


We are badly behind the schedule we proposed in the APP and need to get a package to procurement ASAP. Tim pointed out that we have often started the procurement process without the final version of specs for other contracts.

New Action Items

  1. Look at 8-stage tubes. -> Sasha
  2. Submit a package to procurement for JLab/FSU contract. -> Mark